OpenStudy (anonymous):

Find the exact time it takes a stone to reach the ground if it is dropped from a bridge that is 200 ft. high. --- "The time t it takes an object dropped from a height h to reach the ground is given by the equation t=(square rooted)h/16

OpenStudy (anonymous):

put h = 200ft in your equation and get t.

OpenStudy (anonymous):

I did that but i think i got the wrong answer, so what did you get?

OpenStudy (anonymous):

I got the answer 3.53 but got it wrong, so...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

can you put up your equation more clearly? is it t = sqrt(h/16) or t = sqrt(h)/16

OpenStudy (anonymous):

t=sqrt(h/16)

OpenStudy (anonymous):

and h is in feet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it has to be t = sqrt(200/16) = 3.5 seconds

OpenStudy (anonymous):

wow, so i just added the extra 3 at the end and got it marked wrongg...

OpenStudy (anonymous):

or actually 3.5 unit time measure.

OpenStudy (anonymous):

yeah i said 3.53 seconds

OpenStudy (anonymous):

the correct answer is 3.53553391 unit time measure.

OpenStudy (anonymous):

sigh. no one actually would say that though , some "real life problem'. i dislike my teacher.

OpenStudy (anonymous):

you shouldn't have said seconds. you probably got marked wrong since you answered in seconds. ask your teacher why they marked it wrong. Because your answer is correct.

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

unless the problem specified that t was in seconds, you shouldn't have answered in seconds.